[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Oct 19, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Wed Oct 19 14:27:12 UTC 2016


Here is today's summary of NHC computer operations:
--- Mello

----------

TAFB:
Update on cumulus performance issues. Mark is working with NCO to find the  
proper video driver and card. The problem might persist for a day or two.

Mello extended the new rwps out to 156 hours for both the Pacific and  
Atlantic domains in GFE.

The long term problem launching the clipping scripts on Crappie is related  
to the menus. The scripts run fine from the command line. TSB will look  
into this week.

HSU:
The SDM informed NHC that supercomputer "tide" is down for maintenance.  
Unfortunately, network administrators did not configure the backup computer  
network settings properly. Until this is fixed, forecasters need to take  
additional steps according to Monica.
If the ATCF hangs after submitting guidance and (terminal window asking for  
password) hit Control C twice; this will get the ATCF past that portion of  
the script that causes the hang-ups. This should allow guidance to be  
submitted. Forecasters indicated no problems with the retrieval process.  
This may need to be done at 12z until the issuance is resolve. These  
instructions have been left on the ATCF.
Other items:

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161019/f5e1268d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list